# thanks: Corresponding author

### v1

@@ Abstract
We test Deep-Learning (DL) techniques for the analysis of rotational core-collapse supernovae (CCSN) gravitational-wave (GW) signals by performing classification and parameter inference of the GW strain amplitude ($D\cdot\Delta h$) and the maximum (peak) frequency $(f_{\textrm{peak}})$, attained at core bounce. Our datasets are built from a catalog of numerically generated CCSN waveforms assembled by Richers et al. 2017. Those waveforms are injected into noise from the Advanced LIGO and Advanced Virgo detectors corresponding to the O2 and O3a observing runs. For a signal-to-noise ratio (SNR) above 5, our classification network using time series detects Galactic CCSN GW signals buried in detector noise with a false positive rate (FPR) of 0.10% and a 98% accuracy, being able to detect all signals with SNR > 10. The inference of $f_{\textrm{peak}}$ is more accurate than for $D\cdot\Delta h$, particularly for our datasets with the shortest time window (0.25 s) and for a minimum SNR = 15. From the calibration plots of predicted versus true values of the two parameters, the standard deviation ($\sigma$) and the slope deviation with respect to the ideal value are computed. We find $\sigma_{D\cdot\Delta h}=52.6$ cm and $\sigma_{f_{\textrm{peak}}}=18.3$ Hz, with respective slope deviations of 11.6% and 8.3%. Our best model is also tested on waveforms from a recent CCSN catalog built by Mitra et al. 2023, different from the one used for the training. For these new waveforms the true values of the two parameters are mostly within the $1\sigma$ band around the network’s predicted values. Our results show that DL techniques hold promise to infer physical parameters of Galactic rotational CCSN even in the presence of real (non-Gaussian) noise conditions from current GW detectors.
@@ I Introduction
The gravitational collapse of the core of massive stars and the subsequent explosion as a supernova (a core-collapse supernova event, or CCSN hereafter) is one of the most interesting sources of gravitational waves (GW) to be detected in the coming years (see [1] for a recent review). CCSN are thought to be the end result of stars with a zero age main sequence (ZAMS) mass larger than about $8M_{\odot}$. The core of these stars, consisting of iron-group nuclei, can no longer sustain nuclear burning and collapses. As the core reaches densities above nuclear saturation density and the equation of state (EoS) stiffens, the infalling material bounces back launching an outward-moving shock wave. Numerical simulations have long helped understand the processes by which the shock wave can gain sufficient energy to power a supernova explosion. Two main mechanisms have been proposed, neutrino-driven explosions and magneto-rotational explosions (see [2, 1] and references therein). Rotation strongly affects the dynamics of CCSN and, in turn, its GW emission. Most CCSN are non-rotating or slowly rotating [3]. In these models, the GW emission is largely stochastic as it is triggered by convection and by the standing accretion shock instability. In rapidly rotating stars, however, the early part of the GW signal is well determined. In these models the explosion is powered by the rotational kinetic energy (see [4] and references therein), leading to a bounce GW signal that is well understood. The GW strain of rapidly-rotating cores reaches its maximum right before bounce and its value depends on the corresponding degree of oblateness of the core. Numerical simulations have shown that for about 10 ms after core bounce the newly formed proto-neutron star (PNS) undergoes a series of oscillations driven by the excitation of axisymmetric fluid modes. Shortly after the GW signal becomes stochastic as a result of the growth of non-axisymmteric hydrodynamical instabilities. One to three CCSN events are expected to occur per galaxy per century [5, 6] and less than 10% are likely to be from fast-rotating progenitors [7].
@@
To date, all confident GW signals observed by the Advanced LIGO and Advanced Virgo detectors correspond to coalescing compact binaries (CBCs) [8, 9, 10, 11]. The most recent optically-targeted searches for GW from CCSN, using data from the first and second observing runs of LIGO and Virgo, found no evidence of significant candidates [5, 12]. As the sensitivity of the detectors increases the probability of detecting GW signals from CCSN will also increase, especially within our Galaxy and nearby satellite galaxies [5, 13, 14, 12, 15, 16, 17]. The successful detection of GW signals from CCSN may help clarify the underlying physical processes occurring in the cores of massive stars during their gravitational collapse along with providing a brand new channel to infer properties of the source, allowing to probe the properties of the PNS, the nuclear EoS, the rotation of the core, and the explosion mechanism [18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 17, 29, 30, 31, 32].
@@
The detection of GW signals from CBCs is achieved through matched filtering [33]. This requires template waveform models to be pre-computed with a faithful representation of true GW. On the contrary, by their stochastic nature, CCSN signals are unmodelled, requiring a completely different approach for their detection. In this case, a coherent time-frequency analysis of the data in a network of detectors is used [34, 12]. Correspondingly, the estimation of the source parameters in the case of CBC signals is done through Bayesian inference. This method uses a large number of waveforms (or “approximants”) that cover a wide parameter space and applies stochastic sampling techniques to evaluate their likelihood functions. Besides being computationally expensive, it may become unsustainable as the number of detections increases (but see [35, 36] for likelihood-free alternatives, and [37, 38] for novel computational proposals). However, there are various reasons why in general a Bayesian approach is not the optimal one for inference of CCSN GW. On the one hand, post-bounce signals rapidly become stochastic a few milliseconds after core bounce. On the other hand, the amount of CCSN waveforms available is severely limited by the computational cost of the simulations and by the impossibility to build the waveforms using the same standard approximations employed for compact binaries. For the fairly brief, “deterministic” part of the early bounce signal, matched-filtering analysis or Bayesian model selection can be used to infer source properties (see e.g. [18, 29]). We also note that the dominant features of CCSN waveforms can be extracted with Principal Component Analysis (PCA), where a mapping between their measured eigenvectors and the physical parameters of the progenitor star can be created [24]. Moreover, for close enough sources, it is also possible to distinguish neutrino-driven CCSN from magneto-rotational CCSN using PCA [39, 40, 19, 41, 30].
@@
An alternative to the methods mentioned before are Machine Learning (ML) and Deep Learning (DL) techniques which have been shown to be significantly less time-consuming than Bayesian approaches in performing parameter estimation of CBC signals [42, 35, 43, 44, 36, 45, 46]. The potential of ML for GW data analysis is becoming increasingly relevant and the number and scope of applications is already important (see [47, 48, 49, 50] for recent reviews). Those include, e.g., detection methods for CBC signals [51, 52, 53], signal quality improvements [54, 55], waveform generation [56, 57, 58], and noise transients simulations in detectors [59, 60]. In the context of CCSN GW signals the application of ML techniques is also an active field of research [61, 62, 63, 28, 64, 32, 30, 65]. The potential of Convolutional Neural Networks (CNNs) to detect CCSN GW signals was first shown by [61] using phenomenological waveforms injected in Gaussian noise. CNNs were also employed by [62] with CCSN GW signals injected in detector noise. At a false alarm probability of 10%, neutrino-driven explosions at 10 kpc yield an expected true alarm probability of 55% (76%) for current (future) LIGO-Virgo-KAGRA (LVK) detectors, while for magneto-rotational explosions at 50 kpc the corresponding values increase to 84% (92%). Approaches based on learned dictionaries were first applied in [28] for CCSN signal classification, obtaining 85% true classifications on signals with a SNR from 15 to 20 (see also [30]). In [63] a Mini-Inception Residual Network (ResNet) for time-frequency images (also implementing convolutional layers) was used to improve the detectability of CCSN GW signals with the LVK pipeline coherent Wave Burst [34]. For a dataset comprising phenomenological CCSN waveforms in background noise from O2, an efficiency of around 80% was obtained for ${\rm SNR}\sim 16$. Recently [32] have used ML to investigate if it is possible to infer the iron core mass from the bounce and early ring-down GW signal of rapidly rotating CCSN models and [64] have compared the performance of different CNNs and long short-term memory networks for multi-label classification of CCSN simulated signals and noise transients using real data. CNNs have also been used by [65] to classify GW from CCSN using spectrograms from numerical simulations injected onto real noise data from O3. We also note the recent work by [30] which presents a comprehensive comparison of different methods (Bayesian model selection, dictionary learning, and CNNs) to determine the explosion mechanism from a GW CCSN detection using up-to-date waveforms from simulations (including, in particular, new 3D long-duration magneto-rotational CCSN waveforms that cover the full explosion phase).
@@
In this work, we present a DL approach for parameter estimation using GW from rapidly rotating CCSN. We train Residual CNNs using datasets built from the numerically generated CCSN waveforms computed by [20] which are injected into noise from the O2 and O3a runs of the Advanced LIGO and Advanced Virgo detectors. In addition, our best model is also tested on a different set of waveforms for rapidly rotating CCSN built by [32] which allows us to assess the generality of our approach. We follow closely the recent work of [29] and perform our inference on two key parameters, the GW strain amplitude, $D\cdot\Delta h$, and the maximum (peak) frequency, $f_{\textrm{peak}}$, attained at core bounce. As discussed in [29], the results from [20] for a 12 $M_{\odot}$ progenitor and a large group of EoS show that when the ratio of rotational-kinetic energy to gravitational potential energy, $T/|W|$, is below 0.06 (i.e. for slowly rotating cores), the first parameter is proportional to $T/|W|$ and the second one is proportional to the square root of the central density, $\sqrt{\rho_{c}}$. These relationships allow, in principle, the inference of PNS properties from the GW detection of a rotating CCSN. In [29] inference on these two parameters was conducted using Bayesian model selection employing a master waveform template built from the waveform catalog of [20]. Our approach, based on DL, can thus be regarded as complementary to the Bayesian approach of [29]. As we show in this work, for a rotating Galactic CCSN, the inference of these two parameters using neural networks can be successfully achieved, yielding results within the $1\sigma$ band from the expected (true) values.
@@
This paper is organized as follows: In Section II we describe the CCSN waveforms we select to train our models. In Section III we discuss a first approach in which the analysis of the signals is performed using spectrograms, focusing on signal classification only. Section IV presents a second, more complete approach. Here, DL methods are applied to time series and we discuss results for both signal classification and parameter inference. Finally, our conclusions are outlined in Section V.
@@ II Waveform selection
The waveforms were obtained through a combination of 1D simulations of the collapse-phase deleptonization with the code GR1D [72] and 2D core-collapse simulations with the CoCoNuT Code [73]. To produce the collapse-phase deleptonization, Richers et al. used an approximation proposed by Liebendörfer [74] where the electron fraction ($Y_{e}$) is parameterized as a function of density ($\rho$) only from spherically symmetric (1D) non-rotating general relativity hydrodynamic simulations. This approximation was also used for the rotating case, as electron captures and neutrino interactions with matter are local and depend on the density in the collapse phase. In this case, the rotational flattening of the collapsing core can be considered relatively small. The axisymmetric (2D) CCSN simulations were performed assuming the Conformal Flat Condition (CFC) approximation for the spacetime metric and an initial rotational profile ($\Omega$) of the pre-collapse star given by
$\Omega(\varpi)=\Omega_{0}\left[1+\left(\frac{\varpi}{A}\right)^{2}\right]^{-1},$ (1)
where $\Omega_{0}$ is the maximum initial (pre-collapse) rotation rate, $\varpi$ is the distance from the axis of rotation and $A$ is a measure of the degree of differential rotation.
@@
The waveforms assume an optimally oriented source. Therefore, for an axisymmetric system there is no cross-polarization ($h_{\times}$) and only the plus-polarization component ($h_{+}$) is different from zero, i.e. $h_{\times}=0$ and $\quad h_{+}=h_{+}^{\rm opt}\sin^{2}\theta$. Here, $\theta$ is the inclination angle between the rotational axis of the core and the line of sight of the observer. We fix this angle to $\pi/2$ rad, simplifying the expression above to $h_{+}=h_{+}^{\textrm{opt}}$. Each waveform on the catalog provides several physical parameters, from which we used the strain in units of distance ($D\cdot h_{+}^{\textrm{opt}}$), the timestamps $(t-t_{\textrm{b}})$, the peak frequency ($f_{\textrm{peak}}$), the maximum rotational velocity $\Omega_{0}$, and the strain amplitude ($D\cdot\Delta h$).
@@ III.1 Dataset generation and neural network details
Our first analysis of the waveforms is based on time-frequency representations of the signals, or spectrograms. These diagrams represent the signal amplitude in pixel intensity, with time on the $x$-axis and frequency on the $y$-axis. The use of spectrograms is not arbitrary. Since the bounce and post-bounce phases of CCSN involve many frequencies, more significant patterns may be found in the time-frequency representation (produced by splitting the input time series into shorter sections and taking Fourier/wavelet transforms of these sections) than in the time-strain representation. This pattern, being contained in a short time interval and having many frequencies, is seen in a spectrogram as an intense narrow vertical band, as shown in Figure 1 (right).
@@
In order to train and validate a neural network (NN) it is necessary to have a large enough data bank. For this purpose, we use a dataset generated with all the 1824 waveforms from the Richers et al. catalog [20], as described in Section II. Each of these waveforms is injected at a random time position into sections of actual detector noise from O2 (corresponding to August 2017), with a duration of 4096 ms. We assume a constant sky position, fixing the distance to the source to 20 kpc and the right ascension, declination, and polarization angle to zero. As the detectors do not have the same sensitivity in the whole frequency range, a whitening procedure is applied to normalize the intensity of each frequency. The frequency range we consider is between 20 and 1000 Hz, where the LIGO and Virgo detectors are most sensitive. We also notch out fixed frequencies affected by known artifacts, namely the frequencies at 60, 120, and 240 Hz for the LIGO detectors and at 50, 100, and 200 Hz for the Virgo detector, corresponding to the US and European power grid frequency and their first harmonics, respectively. The sampling frequency is set to 16384 Hz. High-resolution time-frequency maps of the injected waveforms are constructed using the Q-transform [75].
@@
The resulting images have dimensions of 256$\times$256 pixels. Since we use three detectors, the images are coded with RGB colors where red, green, and blue correspond to the spectrograms of LIGO-Hanford, LIGO-Livingston, and Virgo detectors, respectively. In addition, the intensity of each color varies between 0 and 255, so that when injecting a very intense signal, the normalization makes the noise almost imperceptible. In the left panel of Figure 1 an RGB image of only real background noise is shown, while the right panel displays an RGB image with an injected CCSN GW signal. In the latter, the signal stands out over the LIGO-Hanford and LIGO-Livingston noise. It shows two narrow stripes with a rapid variation in frequency, associated with the burst, and two subsequent features appearing shortly after, connected with the oscillations of the PNS. The noise is barely noticeable in the green and red channels due to the normalization while in the blue channel the intensity of the injections is much lower than the noise (hence, the normalization makes the noise stand out). Moreover, a delay can be observed between the detection of the two interferometers.
@@
A bank of $10^{4}$ images was generated, where half of them are only noise (labeled as “background”) and the other half have CCSN GW signals injected into noise (labeled as “signal”). In addition, the SNR of each signal was evaluated, serving as a filter to eliminate weak signals (SNR < 5) that would otherwise only confuse our NN. Subsequently, and for a better understanding of our models, two other sets of images were generated for testing. One dataset consists of 1500 images arranged in groups differentiated by five SNR units, and the other dataset consists of 4000 images with distances ranging between 5 kpc and 40 kpc.
The SNR of a signal is a combination of the individual SNR for each GW interferometer in the network, LIGO-Hanford, LIGO-Livingston and Virgo, defined as
@@
The individual $\textrm{SNR}_{i}$ is found by calculating the cross-correlation matrix between the post-injected strain and the injected template after projecting into the corresponding detector. Both strains are treated with whitening, band-pass and notch filtering, as explained before. The maximum value of this cross-correlation matrix is considered as the $\textrm{SNR}_{i}$.
As our study is based on images we work with CNNs. These must be capable of differentiating between background and signal. We use a pre-trained ResNet101 model which implements an architecture with a residual CNN 101 layers deep [76]. We load the ResNet101 model pre-trained with over one million images from the ImageNet database [77]. The pre-trained network can classify images into $1000$ object categories. As a result, the network has learned feature-rich representations for a wide range of images, although we will only need it to distinguish between two classes, background and signal. To accomplish this, we need to adapt the network by changing its last layer of 1000 neurons to one with only 2 of them. For the generation of the datasets, functions from the python [78] libraries SciPy [79], PyCBC [80], and GWpy [75], were used. We also used the libraries fast.ai [81] for the classification tests.

### v2

@@ Abstract
We test deep-learning (DL) techniques for the analysis of rotational core-collapse supernovae (CCSN) gravitational-wave (GW) signals by performing classification and parameter inference of the maximum (peak) frequency and the GW strain amplitude ($\Delta h$) multiplied by the luminosity distance ($D$) attained at core bounce, respectively, $(f_{\textrm{peak}})$ and $(D\cdot\Delta h)$. Our datasets are built from a catalog of numerically generated CCSN waveforms assembled by Richers et al. 2017. Those waveforms are injected into noise from the Advanced Laser Interferometer Gravitational Wave Observatory and Advanced Virgo detectors corresponding to the O2 and O3a observing runs. For a network signal-to-noise ratio (SNR) above 5, our classification network using time series detects Galactic CCSN GW signals buried in detector noise with a false positive rate of 0.10% and a 98% accuracy, being able to detect all signals with SNR > 10. The inference of $f_{\textrm{peak}}$ is more accurate than for $D\cdot\Delta h$, particularly for our datasets with the shortest time window (0.25 s) and for a minimum SNR = 15. From the calibration plots of predicted versus true values of the two parameters, the standard deviation ($\sigma$) and the slope deviation with respect to the ideal value are computed. We find $\sigma_{D\cdot\Delta h}=52.6$ cm and $\sigma_{f_{\textrm{peak}}}=18.3$ Hz, with respective slope deviations of 11.6% and 8.3%. Our best model is also tested on waveforms from a recent CCSN catalog built by Mitra et al. 2023, different from the one used for the training. For these new waveforms, the true values of the two parameters are mostly within the $1\sigma$ band around the network’s predicted values. Our results show that DL techniques hold promise to infer physical parameters of Galactic rotational CCSN events.
@@ I Introduction
The gravitational collapse of the core of massive stars and the subsequent explosion as a supernova (a core-collapse supernova event, or CCSN event hereafter) is one of the most interesting sources of gravitational waves (GWs) to be detected in the coming years (see [1] for a recent review). CCSN events are thought to be the end result of stars with a zero age main sequence mass larger than about $8M_{\odot}$. The core of these stars, consisting of iron-group nuclei, can no longer sustain nuclear burning and collapses. As the core reaches densities above nuclear saturation density and the equation of state (EOS) stiffens, the infalling material bounces back launching an outward-moving shock wave. Numerical simulations have long helped understand the processes by which the shock wave can gain sufficient energy to power a supernova explosion. Two main mechanisms have been proposed, neutrino-driven explosions and magneto-rotational explosions (see [2, 1] and references therein). Rotation strongly affects the dynamics of CCSN events and, in turn, its GW emission. Most CCSN events are nonrotating or slowly rotating [3]. In these models, the GW emission is largely stochastic as it is triggered by convection and by the standing accretion shock instability. In rapidly rotating stars, however, the early part of the GW signal is well determined. In these models the explosion is powered by the rotational kinetic energy (see [4] and references therein), leading to a bounce GW signal that is well understood. The GW strain of rapidly-rotating cores reaches its maximum right before bounce and its value depends on the corresponding degree of oblateness of the core. Numerical simulations have shown that for about 10 ms after core bounce the newly formed proto-neutron star (PNS) undergoes a series of oscillations driven by the excitation of axisymmetric fluid modes. Shortly after the GW signal becomes stochastic as a result of the growth of nonaxisymmetric hydrodynamical instabilities. One to three CCSN events are expected to occur in the Milky Way per century [5, 6] and less than 10% are likely to be from fast-rotating progenitors [7].
@@
To date, all confident GW signals observed by the Advanced Laser Interferometer Gravitational Wave Observatory (LIGO) and Advanced Virgo detectors correspond to coalescing compact binaries (CBCs) [8, 9, 10, 11]. The most recent optically-targeted searches for GWs from CCSN events, using data from the first three observing runs of LIGO and Virgo, found no evidence of significant candidates [5, 12]. As the sensitivity of the detectors increases the probability of detecting GW signals from CCSN events will also increase, especially within our Galaxy and nearby satellite galaxies [5, 13, 14, 12, 15, 16, 17]. The successful detection of GW signals from CCSN events may help clarify the underlying physical processes occurring in the cores of massive stars during their gravitational collapse along with providing a brand new channel to infer properties of the source, allowing to probe the properties of the PNS, the nuclear EOS, the rotation of the core, and the explosion mechanism [18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 17, 29, 30, 31, 32].
@@
The detection of GW signals from CBCs is achieved through matched filtering [33]. This requires template waveform models to be precomputed with a faithful representation of true GWs. On the contrary, by their stochastic nature, CCSN GW signals are unmodeled, requiring a completely different approach for their detection. In this case, a coherent time-frequency analysis of the data in a network of detectors is used [34, 12]. Correspondingly, the estimation of the source parameters in the case of CBC signals is done through Bayesian inference. This method uses a large number of waveforms (or “approximants”) that cover a wide parameter space and applies stochastic sampling techniques to evaluate their likelihood functions. Being computationally expensive, the Bayesian approach is, in general, not the optimal method to perform parameter inference of CCSN GW signals for several reasons. On the one hand, postbounce signals rapidly become stochastic a few milliseconds after core bounce. On the other hand, the amount of CCSN waveforms available is severely limited by the computational cost of the simulations and by the impossibility to build the waveforms using the same standard approximations employed for compact binaries. For the fairly brief “deterministic” part of the early bounce signal, matched-filtering analysis or Bayesian model selection can be used to infer source properties (see, e.g., [18, 29]). We also note that the dominant features of CCSN waveforms can be extracted with principal component analysis (PCA), where a mapping between their measured eigenvectors and the physical parameters of the progenitor star can be created for third-generation detectors [24]. Moreover, for close enough sources, it is also possible to distinguish neutrino-driven CCSN events from magneto-rotational CCSN events using PCA [35, 36, 19, 37, 30].
@@
An alternative to the methods mentioned before are machine learning (ML) techniques which have been shown to be significantly less time-consuming than Bayesian approaches in performing parameter estimation of CBC signals [38, 39, 40, 41, 42, 43, 44]. The potential of ML for GW data analysis is becoming increasingly relevant, and the number and scope of applications are already important (see [45, 46, 47, 48] for recent reviews). Those include, e.g., detection methods for CBC signals [49, 50, 51], signal quality improvements [52, 53], waveform generation [54, 55, 56], and noise transient simulations in detectors [57, 58]. In the context of CCSN GW signals the application of ML techniques is also an active field of research [59, 60, 61, 28, 62, 32, 30, 63]. The potential of convolutional neural networks (CNNs) to detect CCSN GW signals was first shown by [59] using phenomenological waveforms injected in Gaussian noise. CNNs were also employed by [60] with CCSN GW signals injected in detector noise. At a false alarm probability of 10%, neutrino-driven explosions at 10 kpc yield an expected true alarm probability of 55% (76%) for current (future) LIGO-Virgo-KAGRA (LVK) detectors, while for magneto-rotational explosions at 50 kpc the corresponding values increase to 84% (92%). Approaches based on learned dictionaries were first applied in [28] for CCSN signal classification, obtaining 85% true classifications on signals with a SNR from 15 to 20 (see also [30]). In [61] a mini-inception residual network (ResNet) for time-frequency images (also implementing convolutional layers) was used to improve the detectability of CCSN GW signals with the LVK pipeline coherent wave burst [34]. For a dataset comprising phenomenological CCSN waveforms in background noise from O2, an efficiency of around 80% was obtained for ${\rm SNR}\sim 16$. Recently [32] have used ML to investigate if it is possible to infer the iron core mass from the bounce and early ring-down GW signal of rapidly rotating CCSN models, and [62] have compared the performance of different CNNs and long short-term memory networks for multilabel classification of CCSN simulated signals and noise transients using real data. CNNs have also been used by [63] to classify GW from CCSN events using spectrograms from numerical simulations injected onto real noise data from O3. We also note the recent work by [30] which presents a comprehensive comparison of different methods (Bayesian model selection, dictionary learning, and CNNs) to determine the explosion mechanism from a GW CCSN detection using up-to-date waveforms from simulations (including, in particular, new three-dimensional (3D) long-duration magneto-rotational CCSN waveforms that cover the full explosion phase).
@@
In this work, we present a deep learning (DL) approach for classification and parameter estimation using GWs from rapidly rotating CCSN events. We test residual CNN algorithms [64, 65] using datasets built from the numerically generated CCSN waveforms computed by [20], which are injected into noise from the Advanced LIGO and Advanced Virgo detectors. We initially test these algorithms with spectrograms for classification as a proof of concept, building on previous group results on CBCs [49]. Additionally, we also test the same algorithms for classification and parameter estimation using time series. A different set of waveforms for rapidly rotating CCSN events, developed by [32], is used to assess the generality of our approach. The tests in the time domain follow closely the recent work of [29] and perform our inference on two key parameters, the GW strain amplitude ($\Delta h$) multiplied by the luminosity distance ($D$), $D\cdot\Delta h$, and the maximum (peak) frequency, $f_{\textrm{peak}}$, attained at core bounce. As discussed in [29], the results from [20] for a 12 $M_{\odot}$ progenitor and a large group of equations of state show that when the ratio of rotational-kinetic energy to gravitational potential energy, $T/|W|$, is below 0.06 (i.e. for slowly rotating cores), the first parameter is proportional to $T/|W|$, and the second one is proportional to the square root of the central density, $\sqrt{\rho_{c}}$. These relationships allow, in principle, the inference of PNS properties from the GW detection of a rotating CCSN event. In [29] inference on these two parameters was conducted using Bayesian model selection employing a master waveform template built from the waveform catalog of [20]. Our approach, based on DL, can thus be regarded as complementary to the Bayesian approach of [29]. As we show in this work, for rotating Galactic CCSN events, the inference of these two parameters using neural networks can be successfully achieved, yielding results within the $1\sigma$ band from the expected (true) values.
@@
This paper is organized as follows: In Sec. II we describe the CCSN waveforms we select to train our models. In Sec. III we discuss a first method in which the analysis of the signals is performed using spectrograms, focusing on signal classification only. Section IV presents a second, more complete method. Here, DL methods are applied to time series and we discuss results for both signal classification and parameter inference. Finally, our conclusions are outlined in Sec. V.
@@ II Waveform selection
The waveforms were obtained through a combination of 1D simulations of the collapse-phase deleptonization with the code GR1D [72] and 2D core-collapse simulations with the CoCoNuT Code [73]. To produce the collapse-phase deleptonization, Richers et al. used an approximation proposed by Liebendörfer [74] where the electron fraction ($Y_{e}$) is parametrized as a function of density ($\rho$) only from spherically symmetric (1D) nonrotating general relativity hydrodynamic simulations. This approximation was also used for the rotating case, as electron captures and neutrino interactions with matter are local and depend on the density in the collapse phase. In this case, the rotational flattening of the collapsing core can be considered relatively small. The axisymmetric (2D) CCSN simulations were performed assuming the conformal flat condition approximation for the spacetime metric, and by forcing the initial model to rotate with constant angular velocity according to the rotation law
$\Omega(r)=\Omega_{0}\left[1+\left(\frac{r}{A}\right)^{2}\right]^{-1},$ (1)
where $\Omega(r)$ is the angular velocity, $r$ is the distance from the axis of rotation, and $\Omega_{0}$ and $A$ are free parameters that determine the rotational speed/energy of the model and the distribution of angular momentum.
@@
The waveforms assume an optimally oriented source. Therefore, for an axisymmetric system there is no cross-polarization ($h_{\times}$), and only the plus-polarization component ($h_{+}$) is different from zero, i.e., $h_{\times}=0$ and $\quad h_{+}=h_{+}^{\rm opt}\sin^{2}\theta$. Here, $\theta$ is the inclination angle between the rotational axis of the core and the line of sight of the observer. We fix this angle to $\pi/2$ rad, simplifying the expression above to $h_{+}=h_{+}^{\textrm{opt}}$. Each waveform on the catalog provides several physical parameters, from which we used the strain in units of distance ($D\cdot h_{+}^{\textrm{opt}}$), the timestamps $(t-t_{\textrm{b}})$, the peak frequency ($f_{\textrm{peak}}$), the maximum rotational velocity $\Omega_{0}$, and the GW strain amplitude ($\Delta h$) multiplied by the luminosity distance ($D$), ($D\cdot\Delta h$).
@@ III.1 Dataset generation and neural network details
Our first analysis of the waveforms is based on time-frequency representations of the signals or spectrograms. These diagrams represent the signal amplitude in pixel intensity, with time on the $x$ axis and frequency on the $y$ axis. The use of spectrograms is not arbitrary. Since the bounce and postbounce phases of CCSN events involve many frequencies, different patterns may be found in the time-frequency representation (produced by splitting the input time series into shorter sections and taking Fourier/wavelet transforms of these sections) when compared with the time-strain representation. These patterns, being contained in a short time interval and having many frequencies, are seen in a spectrogram as an intense narrow vertical band, as shown in Fig. 1 (right).
@@
In order to train and validate a neural network (NN) it is necessary to have a large enough data bank. To perform the classification test, we created a dataset consisting of images of only noise and others with injected signals. For both signal and noise spectrograms, we randomly selected time series segments of actual detector noise from O2 (corresponding to August 2017), with a duration of 4096 ms, without any specific criteria. We normalized the intensity of each frequency using a whitening procedure to account for variations in detector sensitivity across the frequency range. Our analysis considers frequencies between 20 and 1000 Hz, where the LIGO and Virgo detectors are most sensitive. We also notched out fixed frequencies affected by known artifacts, specifically 60, 120, and 240 Hz for the LIGO detectors and 50, 100, and 200 Hz for the Virgo detector, corresponding to the U.S. and European power grid frequency and their first harmonics, respectively. The sampling frequency was set to 16384 Hz. High-resolution time-frequency maps with a final time duration of 200 ms were constructed using the Q-transform method from GWpy [75], with its default parameters. For the signal spectrograms, the only difference was the injection of a waveform at a random time position into a section of actual detector noise before whitening. We utilize all the 1824 waveforms from the Richers et al. catalog [20], described in Sec. II, assuming a constant sky position, fixing the distance to the source to 20 kpc and the right ascension, declination, and polarization angle to zero.
@@
The resulting images have dimensions of 256$\times$256 pixels, where the frequencies and time are distributed linearly on their respective axis. Since we use three detectors, the images are coded with RGB colors where red, green, and blue correspond to the spectrograms of LIGO-Hanford, LIGO-Livingston, and Virgo detectors, respectively. In addition, the intensity of each color varies between 0 and 255, so that when injecting a very intense signal, the normalization makes the noise almost imperceptible. In the left panel of Fig. 1 a RGB image of only real background noise is shown, while the right panel displays a RGB image with an injected CCSN GW signal. In the latter, the signal stands out over the LIGO-Hanford and LIGO-Livingston noise. It shows two narrow stripes with a rapid variation in frequency, associated with the burst, and two subsequent features appearing shortly after, connected with the oscillations of the PNS. The noise is barely noticeable in the green and red channels, due to the normalization, while in the blue channel, the intensity of the injections is much lower than the noise (hence, the normalization makes the noise stand out). Moreover, a delay can be observed between the detection of the two interferometers.
@@
A bank of $10^{4}$ images was generated, where half of them are only noise (labeled as “background”) and the other half have CCSN GW signals injected into noise (labeled as “signal”). This dataset was divided into two: a training set (85%) and a validation set (15%). In addition, the SNR of each signal was evaluated, serving as a filter to eliminate weak signals (SNR < 5) that would otherwise only confuse our NN. We always consider the network SNR amplitude of the signals, which combines the individual SNR for each GW interferometer in the network, LIGO-Hanford, LIGO-Livingston, and Virgo, defined as
@@
The individual $\textrm{SNR}_{i}$ is found by calculating the cross-correlation matrix between the postinjected strain and the injected template after projecting into the corresponding detector. Both strains are treated with whitening, bandpass, and notch filtering, as explained before. The maximum value of this cross-correlation matrix is considered as the $\textrm{SNR}_{i}$.
As our study is based on images, we work with CNNs. These must be capable of differentiating between background and signal. We use a pretrained ResNet101 model which implements an architecture with a residual CNN 101 layers deep [76]. We load the ResNet101 model pretrained with over one million images from the ImageNet database [77]. The pretrained network can classify images into $1000$ object categories. As a result, the network has learned feature-rich representations for a wide range of images, although we will only need it to distinguish between two classes, background and signal. To accomplish this, we need to adapt the network by changing its last layer of 1000 neurons to one with only 2 of them. For the generation of the datasets, functions from the PYTHON [78] libraries SciPy [79], PyCBC [80], and GWpy [75] were used. We also used the library fast.ai [81] for the classification tests.